# SATA-Bench: Select All That Apply Benchmark for Multiple Choice Questions

### v2

@@ 3.1 Key Observations
SATA-Bench is challenging and different. All models have a precision greater than $80\%$, but none achieves an EM score above $42\%$. This indicates that while models often select some correct answers, they fail to consistently identify all of them.
@@
In general, proprietary models have higher EM and Precision than their open-source counterparts. Unlike in other benchmarks, there is no single model that dominates across all performance metrics. Some large reasoning models (LRM), such as O3 and Grok 3 Think, have higher EM and Recall than non-reasoning models. Interestingly, larger and more recent models do not necessarily perform better. For example, Claude 3 Sonnet performs better than Claude 3.5 Sonnet V1 and Claude 3 Opus in exact match rate. However, within the Claude family, larger models always have higher precision. For example, Claude 3 Opus has the highest precision among the Claude 3 model family. According to  [4, 13], these results contrast with existing single-choice MCQ LLM performance, such as MMLU [24] and ARC [10], where larger or more recent models tend to show clear improvements.
@@
Models choose too few answers. Nearly all LLMs tend to select fewer answers than required. As an extreme example, Llama 3.1 70B on average, selects one fewer option per question than the correct number. Accordingly, Llama 3.1 70B achieves the highest precision but the lowest exact match (EM). The tendency to under-select increases as the number of correct answers grows (see Figure 11). This behavior negatively impacts the EM rate for questions with many correct choices (see Figure 12). The highest CtAcc is only $48\%$, even the best model predicts the correct number of answers in fewer than half of the questions. We hypothesize that this behavior results from models being primarily trained and evaluated on benchmarks where each question has only one correct answer, making them poorly suited for SATA tasks. Through a t-test, we observe that the mean of the CtDif column is significantly lower than 0, with a p-value of $1.70\times 10^{-6}$, supporting the observation that models consistently under-select answers.
@@
Unselection bias exists. Some models exhibit a tendency to avoid selecting specific labels, even when they are correct. [^2] However, we cannot exclude position effect [56]. When comparing Selection Probability Divergence (SPD) from our benchmark with 1,000 randomly simulated SPDs, Welch’s t-test reveals that LLMs’ SPD is significantly higher than random, with a p-value of $0.0467$. Even the best model in terms of selection bias (Gemini) underperforms on label M, with its recall rate being 6.3% lower than its average recall (Figure 10).
There is no clear winner across datasets. When breaking down the benchmark by its six datasets, different models excel in different domains. Top-performing models vary by dataset, showcasing the diversity of SATA-BENCH. Considering exact match rate (see Figure 13): O3 excels in News and Events classification. DeepSeek R1 leads in Biomedicine, GPT-4.1 performs best in reading comprehension, and Claude 3.7 dominates in toxicity and laws. This highlights the importance of domain-specific evaluation and the breadth of challenges covered by the benchmark.
@@ 3.2 Ablation Studies
We conducted ablation studies to test different strategies for improving model performance. We report the average results across three models selected for diverse profiles in terms of cost, open-source availability, and overall performance. The complete prompts are provided in Appendix H.3.
Improving performance on SATA-Bench is challenging. We tried several approaches to improve performance, but none yielded consistent or significant improvements.
  • Changing the symbol used for each answer choice did not improve the selection bias. We replaced the default option IDs from A/B/C/D to a/b/c/d and 1/2/3/4. While the 1/2/3/4 format achieved slightly better exact match accuracy, it also increased selection bias and reduced precision. Overall, we did not observe performance improvement by changing symbols (see rows 1-3 in Table 3).
  • We provided few-shot examples in the prompt before the test models. However, this strategy did not lead to meaningful improvements in performance (see row 4 in Table 3).
  • Inspired by survey science [46, 39], we instructed the models to examine each option individually. However, the models still selected too few options overall and did not improve performance (see rows 5-6 in Table 3).
Given more information, two approaches do improve performance and can provide additional insights into why the models struggle.
  • Providing the number of correct choices improves performance. To understand how much error is due to the models’ lack of knowledge regarding the number of correct options, we explicitly provided this information in the instruction for each question. This increased the exact match rate by $20.95$ percentage points and reduced the selection bias metric RStd. However,when we instead provided the average number of correct choices across all questions in SATA-Bench, performance declined (see rows 7-8 in Table 3).
  • Converting questions to multi-choice question with one correct answer. For example, consider a question with three correct answers and six incorrect answers: we expanded it into three separate single-choice questions, each with one correct answer and six incorrect answers. We redefined the exact match rate as the percent of all original questions where a model answered all expanded questions correctly. This approach improved performance by $12.53\%$ (see row 9 in Table 3), demonstrating that SATA questions are significantly harder for LLMs than single choice questions.
Both results suggest that while models can often identify individual correct answers, they lack awareness of how many correct answers exist, which contributes to their low performance.
@@
$NOTA$ stop option, $\tau$ confidence threshold
# Initialize the selected option set
@@ 4 Improving Performance on SATA Questions
The experimental results in Section 3 demonstrate that Selection Bias and Count Bias degrade LLM performance on SATA-Bench, and that simple prompting strategies do not lead to significant improvements. This section focuses on improving performance on open-source models, which allows us to leverage token-level logits or probability estimates from the first token prediction.
@@
To address Selection Bias, we draw from prior research on token debiasing methods [9, 56] in the MCQ setting, where selection bias is attributed to the a priori probability mass assigned by the model to specific option IDs. These methods propose various techniques to capture and remove such biases. We hypothesize that similar debiasing techniques can be adapted to mitigate unselection bias in SATA tasks. To address Count Bias, we retrieve the predicted probabilities of option IDs and select options whose probabilities exceed a predefined threshold. However, because SATA-Bench includes a large option set, the probability distribution tends to decay rapidly, with most options receiving near-zero probability mass beyond the first few choices. This makes it challenging to establish a reliable threshold. Converting SATA questions into multiple binary classification problems helps but significantly increases inference cost.
@@
Choice Funnel Algorithm. To improve model performance on SATA problems, we propose a decoding method called Choice Funnel (see Algorithm 1). This approach first adds an auxiliary option “None of the above”. It then selects the option with the highest first debiased token probability and removes it from the option set. This process repeats iteratively until one of two stopping conditions is met: (i) the model selects the “None of the above” option or (ii) the probability of the next selected option falls below a predefined confidence threshold.
The idea of introducing “None of the above" (NOTA) comes from the traditional survey science domain, where options like “I don’t know" (idk) are commonly included to improve the data collected in surveys [45]. Recent research shows that survey design principles can inform LLM development [17] and that LLMs exhibit similar biased response behaviors as humans [9, 16]. In our case NOTA outperforms idk (see ablation study in Appendix M.1).
@@
The intuition behind the second stopping condition comes from our observation of model output probabilities, where the highest token probability tends to be lower at the beginning of iterations, since the model treats multiple options as equally correct. Later in the process, relatively higher probability is assigned to the final remaining correct option in the option set. We also show that Choice Funnel performs best when both stopping conditions are used together (see ablation study in Appendix M.3). Regarding the choice of DebiasingFunction in Algorithm 1, Choice Funnel is flexible and can incorporate any token debiasing method proven effective in MCQ settings. We demonstrate one such debiasing method in Section 4. Additional ablation results on each sub-component of Choice Funnel are provided in Appendix M.2. Finally, the inference cost of Choice Funnel, measured by the number of model forward passes, scales linearly with the number of correct labels rather than the total number of labels. This makes the method particularly efficient when the correct labels represent a small fraction of the option set.
Experimental Setup. We adapted the PriDe algorithm [56] as the token debiasing method in our experiments due to its label-free and computationally efficient implementation. It works by first estimating the model’s prior bias towards specific option ID tokens (e.g., A, B, C) through random permutations of option contents in a small subset of test samples (10% in our experiments). We then use this estimated prior to adjust the prediction distribution on the remaining samples, separating the model’s inherent positional and token biases from its task-specific predictions. Because the original PriDe algorithm was designed for standard single-answer MCQ settings, we modified it to better fit our SATA setting (see Appendix K).
We evaluate the performance of Choice Funnel against three baseline methods that rely on first-token probabilities: (i) First token probability with a fixed threshold, as defined in Section 3 (referred to as first token). (ii) Building on method 1, we apply PriDe debiasing method [56] (referred to as first token debiasing). (iii) Convert each option into an individual binary yes/no question (referred to as yes/no). We expect yes/no to be a strong baseline, as it evaluates each choice independently. In this study, we use basic prompts (see Appendix H) and experiment with 7 LLMs from Table 2 that fall under the Probability Based Retrieval category (more details in Appendix L). For each model, we compute metrics reported in Table 2, and additionally report an InfCost metric to capture the number of model forward passes required for each method.
Key Observations. Choice Funnel consistently outperforms all three baselines across all 7 models in EM, SPD, and CtAcc (see Table 4). Choice Funnel reduces unselection bias and count bias – compared to the first token baseline, Choice Funnel achieves an average 56.16% reduction in SPD and 154.62% improvement in CtAcc, resulting in a 277.48% gain in Exact Match (EM) performance. While reasoning models also show improvements with Choice Funnel, we exclude these from aggregate calculations as their exceptionally low baselines would artificially inflate gains. When compared to our strongest baseline, the yes/no approach, Choice Funnel achieves a substantial 29.87% improvement in EM while reducing model forward passes by 64.48% thanks to its early stopping mechanism, demonstrating efficient inference scalability. Statistical significance testing (t-test) confirms that Choice Funnel significantly outperforms both yes/no and first token debiasing in EM and CtAcc, with a maximum p-value of $0.0079$. While our models’ parameter sizes (7B-14B) limit direct comparison to much larger proprietary models, Choice Funnel’s performance on the phi3-small model still exceeds that of larger models such as Llama-90B and Mistral-Large V2 (see Table 2). This further underscores the effectiveness of our method. Additional ablation studies on the individual components of Choice Funnel are provided in Appendix M.
@@
Table 4: Performance of various models on SATA-Bench using different decoding methods. Choice Funnel achieves generally better performance, effectively reducing selection and count bias compared to three baseline methods. Best values in each column are highlighted in bold. Columns labeled [$\uparrow$] indicate higher-is-better; columns labeled [$\downarrow$] indicate lower-is-better. All numeric values are rounded to two decimal places.
@@ 5 Related Work
SATA Benchmark. Many existing MCQ benchmarks have only one correct answer and thus cannot test LLMs’ ability to select multiple correct choices. On the one hand, existing SATA datasets, such as [31, 30, 3, 27, 8], have more than 30 labels per question to choose from. This makes it impractical for LLMs to identify all correct labels from such large label pools. Other SATA-style datasets test narrow, specialized capabilities, such as emotional understanding [15] or music style understanding [55], which are less emphasized in mainstream LLM benchmarks. Since most existing methods to solve SATA questions require converting questions to a bag-of-words [32], and as a result, most of the above datasets exist only in bag-of-words format, making them unsuitable for evaluating LLMs in our benchmark setting. To our knowledge, there is currently no existing LLM benchmark that consists exclusively of SATA questions.
Selection Bias. Many previous papers have discussed the tendency of LLMs to favor choices based on option order or specific symbols when answering MCQs [23, 53]. However, these papers have primarily focused on single-answer questions. A common approach to reducing selection bias involves calibrating output probabilities using the prior bias of an option ID [56]. However, it remains unclear how to define or compute such priors in SATA questions.
@@ 6 Conclusion
We introduced SATA-Bench, a carefully curated suite of “Select All That Apply" (SATA) questions designed to evaluate LLMs in scenarios where multiple correct answers must be identified. Spanning diverse question types—from reading comprehension to text classification—and covering domains such as law and biomedicine, SATA-Bench presents a comprehensive challenge for current LLMs. Our benchmarking study of both open-source and proprietary models revealed a best exact-matching accuracy of only $41.8\%$, highlighting the difficulty of selecting all the correct options. Our ablation studies indicate that a simple prompting strategy alone cannot boost exact match performance, primarily because LLMs struggle to determine the correct number of answers. This limitation stems from the fact that SATA-style questions are rarely included in LLMs’ benchmark datasets, making LLMs prone to selection and count biases even when they recognize some correct options. To address this limitation, we proposed the Choice Funnel algorithm, which significantly improves exact match performance by systematically guiding the selection process. Our findings highlight the need for more focused research on handling multiple answer tasks, where partial correctness is insufficient. We hope that SATA-Bench and the choice funnel methodology will encourage the development of more robust LLMs capable of handling realistic, multi-answer scenarios, ultimately improving their effectiveness in real-world applications that require identifying all relevant answers.
Impact Statement
The introduction of SATA-Bench marks a crucial advancement in evaluating Large Language Models (LLMs) on “Select All That Apply" (SATA) multiple-choice questions. By addressing a significant gap in existing benchmarks, which predominantly focus on single-answer multiple-choice tasks, SATA-Bench challenges LLMs with real-world scenarios requiring multiple correct responses across domains such as reading comprehension, law, and biomedicine. This benchmark highlights the limitations of current LLMs, which struggle to accurately determine all valid answers, achieving a best-case exact match accuracy of only $41.8\%$.
SATA-Bench ’s impact extends beyond evaluation, as it reveals key biases in LLM decision-making, such as count bias and selection bias, which hinder performance on multi-answer tasks. To address these shortcomings, the development of the Choice Funnel algorithm demonstrates a novel approach to systematically improving LLM selection accuracy, significantly enhancing model performance in SATA tasks.
While the current focus is on knowledge-intensive domains, the potential for expansion into additional fields such as mathematics, coding, and instruction following is vast. SATA-Bench can also be extended to free-text tasks, where the set of correct responses is not explicitly provided, and to other modalities, such as voice and vision. This would further refine LLM capabilities in handling complex, multi-faceted decision-making tasks. By pushing the boundaries of LLM evaluation, SATA-Bench lays the foundation for the next generation of AI systems capable of more nuanced, flexible reasoning in diverse real-world applications.
Limitations
Memorization. While we believe most questions in SATA-Bench have not been seen during pretraining, we cannot fully rule out the possibility that some LLMs may have been exposed, even partially, to the source datasets. We do not have access to the pretraining data of proprietary models and therefore cannot conclusively assess memorization.
Domain Coverage. SATA-Bench spans six diverse domains, including reading comprehension, biomedicine, and law. However, the total number of domains remains limited compared to larger-scale benchmarks such as MMLU, indicating that further expansion is needed for broader generalization.
Text Modality. Our benchmark is text-only. Real-world tasks often require multimodal reasoning (e.g., interpreting charts, images, or audio), which SATA-Bench does not evaluate. We also do not address other data modalities, such as structured tabular data or sensor data, which are common in practical applications.
Label Correctness. Although we perform rigorous human validation and evaluations, we acknowledge that human beings can make mistakes. Some domains, such as biomedicine and law, are inherently complex and may contain subtle ambiguities. Thus, we cannot guarantee perfect correctness of all labels despite triple human annotation and agreement filtering.
Language Limitation. SATA-Bench includes only English-language questions. Evaluating multilingual capabilities or cross-lingual transfer remains a work for the future.

### v3

@@ 3.2 Key Observations
SATA-Bench is challenging and different. 13 models achieve a JI above $70\%$, but none surpass $42\%$ EM. This shows that while models often identify some correct answers, they fail to consistently recover the full set.
@@
Proprietary models generally achieve higher JI and Precision than open-source ones. Unlike other benchmarks, no single model dominates across all metrics. Notably, larger and more recent models do not always perform better. For instance, Claude 3 Sonnet outperforms Claude 3.5 Sonnet and Claude 3 Opus in exact match, though within the Claude family, larger models consistently have higher precision (e.g., Claude 3 Opus has the highest precision among the Claude 3 variants). According to (Anthropic, 2024; DeepSeek-AI & el at, 2024), these results contrast with performance on single-choice MCQ benchmarks such as MMLU (Hendrycks et al., 2021) and ARC (Clark et al., 2018), where larger or newer models typically show clear gains. Large reasoning models (LRMs) are slightly better than their non-reasoning counterparts in JI but failed to reduce selection and count bias. We provide a case study in Appendix Q to investigate LRM’s behavior.
@@
Models choose too few answers. Nearly all LLMs tend to select fewer answers than required. For example, Llama 3.1 70B selects, on average, one fewer option per question than the correct number. Accordingly, it achieves the highest precision but the lowest Jaccard Index (JI). The tendency to under-select increases as the number of correct answers grows (Figure 11), which in turn depresses JI for questions with many correct choices (Figure 12). Even the best model achieves a CtAcc of only 48%, predicting the correct number of answers in fewer than half of the questions. We hypothesize that this behavior stems from models being primarily trained and evaluated on benchmarks with single correct answers, making them poorly suited for SATA tasks. A t-test confirms this under-selection: the mean of CtDif is significantly below 0, with $p=1.70\times 10^{-6}$.
Models speculate a lot. LLMs also over-select, consistently choosing incorrect options, with all models exceeding a 20% FPR. More than 70% of the models predict at least one incorrect choice more often than they produce exact matches, underscoring their speculating behavior. Interestingly, stronger-performing models tend to speculate more: hallucination rate and exact match are positively correlated ($r=0.61$, $p=8\times 10^{-4}$). This dual trend suggests that as models improve in identifying correct answers, they also become more prone to speculation, highlighting the difficulty of disentangling genuine knowledge from overconfidence in LLM predictions.
@@
Unselection bias exists. Some models exhibit a systematic tendency to avoid selecting certain labels, even when they are correct. When comparing Selection Probability Divergence (SPD) from our benchmark with 1,000 randomly simulated SPDs, Welch’s t-test shows that LLMs’ SPD is significantly higher than random ($p=0.0467$). Even the best model in terms of selection bias (Gemini 2.5) underperforms on label M, with its recall rate 6.3% lower than its overall average recall (Figure 10).
@@ 3.3 Ablation Studies
We conducted ablation studies to test different strategies for improving model performance. We report the average results across three models (Llama 3.1 405B, Nova Pro, Claude 3.5 Haiku) selected for diverse profiles in terms of cost, open-source availability, and overall performance. The complete prompts are provided in Appendix H.3.
We tested multiple strategies to improve performance, but none produced consistent or significant gains, suggesting that prompting alone is insufficient for enhancing SATA performance.
  • Changing option symbols. Replacing the default option IDs (A/B/C/D) with a/b/c/d or 1/2/3/4 did not reduce selection bias. While the numeric format slightly improved exact match, it also increased selection bias and reduced precision. Overall, it’s ineffective.(rows 1–3, Table 3).
  • Few-shot prompting. Providing few-shot examples before test questions produced no meaningful improvements (row 4, Table 3).
  • Option-by-option prompting. Inspired by survey methodology Smyth et al. (2006); Pew Research Center (2019), we instructed models to evaluate each option individually. However, models still under-selected and showed no overall improvement (rows 5–6, Table 3).
With additional information, two strategies improved performance and provided insight into why models struggle:
  • Providing the number of correct answers. To assess how much error stems from uncertainty about the number of valid options, we explicitly told models how many correct answers each question contained. This increased exact match by 20.95 points and reduced selection bias (RStd). However, giving only the average number of correct answers across the dataset reduced performance (rows 7–8, Table 3).
  • Decomposing into single-choice tasks. For a question with three correct and six incorrect options, we converted it into three separate single-choice questions (one correct + six incorrect each). We redefined exact match as the proportion of original questions where all expanded items were answered correctly. This raised performance by 12.53% (row 9, Table 3), showing that SATA questions are much harder for LLMs than single-choice ones.
Together, these results suggest that while models can often identify individual correct answers, their lack of awareness of how many answers to select is a key failure mode, highlighting the need for specialized decoding strategies.
@@
NOTANOTA stop option, $\tau$ confidence threshold
1ex# Initialize the selected option set
@@ 4 Improving Performance on SATA Questions
The experimental results in Section 3 demonstrate that speculation bias, unselection bias, and count bias degrade LLM performance on SATA-Bench, highlighting the need for new decoding algorithms. This section focuses on improving performance in open-source models, which allow us to leverage token-level logits which proprietary models do not expose.
@@
To address unselection bias, we can draw from prior research on token debiasing methods (Choi et al., 2024; Zheng et al., 2024) in the MCQ setting, where selection bias is attributed to the a priori probability mass assigned by the model to specific option IDs. These methods propose various techniques to capture and remove such biases. We hypothesize that these techniques can be adapted to mitigate unselection bias in SATA tasks. To address speculation bias, we want to design a mechanism to encourage LLMs to abstain rather than speculate under uncertainty. To address count bias, we can consider retrieving the predicted probabilities of option IDs and select options whose probabilities exceed a predefined threshold. However, because SATA-Bench includes a large option set, the probability distribution decays rapidly, with most options receiving near-zero probability mass beyond the first few choices. This makes it challenging to establish a reliable threshold. Converting SATA questions into multiple binary classification problems helps but significantly increases inference cost.
@@
Choice Funnel Algorithm. With the above consideration, we propose a decoding method called Choice Funnel (Algorithm 1) tailored to solve SATA problems. This approach first adds an auxiliary option “None of the Above,” then selects the option with the highest debiased token probability and removes it from the option set. The process repeats iteratively until one of two stopping conditions is met: (i) the model selects “None of the Above” or (ii) the probability of the next option falls below a predefined confidence threshold.
The addition of an auxiliary option is inspired by recent research that LLMs exhibit biases similar to those observed in human responses (Choi et al., 2024; Eckman et al., 2024a), aiming to reduce LLM speculation. While “I don’t know” (idk) being the most common option used to improve survey data quality (Schuman & Presser, 1996) and have been suggested in recent LLM research (Kalai et al., 2025), NOTA consistently outperforms idk (see ablation study in Appendix M.1).
@@
The intuition behind the second stopping condition comes from our finding that output probabilities correlate with the number of correct options the model considers: the highest token probability tends to be lower at the beginning of iterations, when the model treats multiple options as equally plausible. Later in the process, relatively higher probability is assigned to the final remaining correct option in the set. We also show that Choice Funnel achieves the best performance when both stopping conditions are used together (see ablation study in Appendix M.3).
Regarding the choice of DebiasingFunction in Algorithm 1, Choice Funnel is flexible and can incorporate any token debiasing method proven effective in MCQ settings. We demonstrate one such method in Section 4. See ablation study on each sub-component in Appendix M.2. Finally, the inference cost of Choice Funnel, measured by the number of model forward passes, scales linearly with the number of correct labels rather than the number of total labels. This makes the method especially efficient when correct labels constitute only a small fraction of the option set.
@@
Table 4: Performance of various models on SATA-Bench using different decoding methods. Choice Funnel achieves consistently stronger results, effectively reducing selection and count bias compared to three baseline methods. The best values in each column are shown in bold. Columns labeled $\uparrow$] indicate higher-is-better, while columns labeled $\downarrow$] indicate lower-is-better. All values are rounded to two decimal places.
@@
Experimental Setup. In our experiments we adapted the PriDe algorithm (Zheng et al., 2024) as DebiasingFunction in Algorithm 1 due to its label-free design and computational efficiency. It works by first estimating the model’s prior bias toward specific option ID tokens (e.g., A, B, C) through random permutations of option contents in a small subset of test samples (10% of the data in our experiments). We then use this estimated prior to adjust the prediction distribution on the remaining samples, thereby separating the model’s inherent positional and token biases from its task-specific predictions. Because the original PriDe algorithm was designed for standard single-answer MCQ tasks, we modified it to better fit the SATA setting (see Appendix K).
We evaluate the performance of Choice Funnel against three baseline methods that rely on first-token probabilities: (i) using the first-token probability with a fixed threshold, as defined in Section 3 (referred to as first token); (ii) applying PriDe debiasing on top of the first-token method (Zheng et al., 2024), current best-performed method in terms of speed and accuracy in solving MCQs. (referred to as first token debiasing); and (iii) converting each option into an individual binary yes/no question (referred to as yes/no). Other advanced calibration methods cannot generalize to SATA or require an extensive dataset to fine-tune the model. In this study, we use standardized prompts (Appendix H) and experiment with seven LLMs from Table 2 that fall under the Probability-Based Retrieval category (details in Appendix L). For each model, we compute the metrics reported in Table 2 and additionally report an InfCost metric to capture the number of model forward passes required for each method.
Key Observations. Choice Funnel consistently outperforms all three baselines across all seven models in EM, SPD, and CtAcc (Table 4). Choice Funnel reduces unselection bias, speculation bias, and count bias—compared to the first token baseline, it achieves an average 56.2% reduction in SPD, 36.4% improvement in JI, and 154.6% improvement in CtAcc, and a 277.5% gain in Exact Match (EM) performance. While reasoning models also show improvements with Choice Funnel, we exclude them from aggregate calculations since their exceptionally low baselines would inflate relative gains. Against the strongest baseline, the yes/no approach, Choice Funnel delivers a 29.9% improvement in EM while reducing model forward passes by 64.5% through its early stopping mechanism, demonstrating scalable inference efficiency. t-test confirms that Choice Funnel significantly outperforms both yes/no and first token debiasing on EM and CtAcc, with a maximum p-value of $0.0079$. Although our models’ parameter sizes (7B–14B) limit direct comparison to much larger proprietary systems, Choice Funnel’s performance on the phi3-small model still surpasses that of larger models such as Llama-90B and Mistral-Large V2 (Table 2), underscoring the effectiveness of our method. Each component of Choice Funnel is essential (Appendix M) and it performs well across larger models (Appendix M.4) and black-box settings (Appendix M.2).
@@ 5 Related Work
SATA Benchmark. Most MCQ benchmarks assume a single correct answer and thus cannot evaluate LLMs’ ability to select multiple options. Existing SATA datasets, such as (Lewis et al., 2004; Kowsari et al., 2017; Aly et al., 2019; Katakis et al., 2008; Charte et al., 2015), often include over 30 labels per question, making exhaustive prediction impractical for LLMs. Others target narrow domains, such as emotion analysis (Demszky et al., 2020) or music style understanding (Zhao et al., 2019), with limited relevance to general reasoning. Many of these datasets are also bag-of-words based (Liu et al., 2022), rendering them unsuitable for evaluating natural question–answer reasoning. Prior work in multi-label classification has explored related text categorization problems (Lewis et al., 2004; Aly et al., 2019), but these methods generally assume bag-of-words features and do not capture LLM reasoning dynamics. No existing LLM benchmark consists exclusively of SATA questions.
Selection Bias. Prior studies show that LLMs favor certain options based on order or symbols when answering MCQs (Gupta et al., 2024b; Wei et al., 2024; Gupta et al., 2024a), though these analyses focus on single-answer settings. Calibration methods using option priors have been proposed (Zheng et al., 2024), but their applicability to SATA tasks remains unclear.
Uncertainty and Survey Methodology. Work on uncertainty quantification has been extensive (Tarekegn et al., 2024b), but is generally framed for probabilistic classifiers rather than multi-answer reasoning. In our setting, uncertainty manifests as systematic speculation bias in LLM predictions. Similarly, survey methodology highlights the role of abstention options such as “I don’t know” or “None of the Above” in reducing respondent bias (Eckman et al., 2024b). Choice Funnel builds on these insights by incorporating abstention to mitigate speculation in SATA tasks.
@@ 6 Conclusion
We introduced SATA-Bench, a dataset of over 10K human-validated SATA questions across six domains, and evaluated 32 LLMs. Even the best model achieves only 41.8% exact match accuracy, with failures driven by three systematic biases: unselection, count, and speculation. Although models can often identify individual correct options, our ablation studies show that they lack reliable mechanisms for estimating the correct number of answers. To address these gaps, we proposed Choice Funnel, a decoding algorithm that combines token debiasing, adaptive thresholding, and abstention handling. Choice Funnel improves the exact match by up to 29 points while reducing the inference cost by 64%, demonstrating that targeted decoding strategies can mitigate systematic errors in multi-answer reasoning. SATA-Bench thus provides both a standardized benchmark and a diagnostic platform to analyze the LLM failure modes. We hope it will guide the development of models better suited for real-world applications where partial correctness is insufficient.
Ethics Statement
Intended Use and Benefits. By diagnosing unselection, speculation, and count biases and proposing a mitigation method (Choice Funnel), this work aims to reduce systematic failure modes that could otherwise yield missed or spurious labels in applications such as content moderation, information extraction, or biomedical tagging. The benchmark is released to facilitate open evaluation and comparative analysis.
Data Provenance and Annotators. Sata-Bench is constructed from publicly available textual sources, carefully filtered and human-validated for clarity and difficulty. We leverage Amazon Bedrock Guardrails to identify and remove any questions containing personally identifiable information (PII). We follow all source licenses and usage policies and do not collect new PII. An internal ethics review was conducted prior to conducting any human annotation or validation for this research.
Avoiding Harm. The goal of this work is to identify and reduce potential harms from LLMs when working on multi-answer questions. To mitigate such risks, we: (1) center the work on evaluation to systematically diagnose where harms arise; (2) report detailed statistics; and (3) propose and benchmark a decoding algorithm that explicitly mitigates these biases.
Reproducibility Statement
Dataset. We describe the SATA transformation process in Section $2.1$, the question-filtering pipeline in Section $2.2$, and the dataset characteristics in Section $2.3$. Complete filtering details—including human validation, redundancy checks, and contamination screening—are provided in Appendix B. A detailed dataset description appears in Appendix A. We release three datasets in the Supplementary Materials: (i) the post-validation set sata-bench-raw-v2.json ($\approx 7.98\text{k}$ items); (ii) the single-choice subset sata-bench-single.json ($\approx 1.57\text{k}$ items); and (iii) the human-annotated set sata-bench-v2.json ($\approx 1.5\text{k}$ items). In total, these releases comprise over 10,00010{,}000 examples.
Evaluation. Evaluation details are described in Section 3. Computational resources used for evaluation are listed in Appendix D. Exact model versions are reported in Table $8$. Inference code is provided in sata_eval.py (Supplementary Materials). Human evaluation procedures are documented in Appendix E. All metrics are detailed in Appendix F, with implementations in metric.py. All prompts are documented in Appendix H. Our handling of inference errors is described in Appendix I. To reproduce inference and ablation studies, run bash run.sh.
Choice Funnel. Choice Funnel is described in Section 4. A detailed description of the benchmarked method appears in Appendix K, with code in debiasing.py. The experimental setup is provided in Appendix L. Ablation studies are reported in Appendices M and E. The full Choice Funnel implementation is provided in the choice_funnel directory in the Supplementary Materials.